OpenStudy (anonymous):

The current I in an electrical conductor varies inversely as teh resistance R of the conductor. The current is 8 amperes when the resistance is 910 ohms. What is the current when the resistance is 530 ohms?

OpenStudy (amistre64):

the kellys had this one as well...... or similar :) i = k/r 8 = k/910 k = 8(910) = 7280 i = 7280/r i = 7280/530 = 13.735

OpenStudy (anonymous):

thank you:)

OpenStudy (amistre64):

youre welcome :)